1b Integrate the following expressions.

∫(3π‘₯3βˆ’4π‘₯2+5π‘₯βˆ’3)𝑑π‘₯

∫(3π‘π‘œπ‘ 2πœƒ+5π‘π‘œπ‘ 4πœƒ)π‘‘πœƒ


1
Expert's answer
2021-04-29T18:10:55-0400

1) "∫(3𝑥^3−4𝑥^2+5𝑥−3)𝑑𝑥=3\int x^3dx-4\int x^2dx+5 \int xdx-3\int 1 \ dx"

"\Rightarrow \dfrac{3}{4}x^4-\dfrac{4}{3}x^3 +\dfrac{5}{2}x^2-3x+C"



2) "\int(3cos2\theta+5cos4\theta)d\theta=3\int cos2\theta d\theta+5\int cos4\theta d\theta"


"\Rightarrow \dfrac{3}{2}sin2\theta+\dfrac{5}{4}sin4\theta+C"